# Discovering Lie Groups with Flow Matching

### v1

@@ 4.2 Training and Generation with LieFlow
Training.
Using exponential curves and the flow matching objective, we can now derive the training and sampling algorithms for our method. The training algorithm (Algorithm 1) directly implements the Lie group flow matching objective by constructing training pairs $(x_{0},x_{1})$ where $x_{1}$ is sampled from the data distribution and $x_{0}=gx_{1}$ for a randomly sampled transformation $g\in G$. $x_{0}$ is now a sample from the prior and lives in the group $G$. Unlike standard conditional flow matching, we construct $x_{0}$ from $x_{1}$ using $g$ and learn the flow over the Lie algebra of group $G$. As $G$ is a group, we can compute the target vector field $A$ analytically using the sampled group element $g$. We sample a timestep $0\leq t\leq 1$ and construct a point $x_{t}=\exp(tA)x_{0}$ on the curve. Given $x_{t}$, the model predicts $A$. This means that the target vector field for this sample pair $(x_{0},x_{1})$ is constant over time. Thus the network only needs to learn to predict the Lie algebra element that generates the transformation from $x_{0}$ to $x_{1}$.
@@
8:  Take gradient descent step on
[…]
10:until converged
@@
5:$M=I$
6:for $t=\{0,\frac{1}{T},\frac{2}{T},\dots,1\}$ do
7:  $A_{t}=v^{\theta}(x_{t},t)$
8:  $x_{t+\Delta_{t}}=\exp(\Delta_{t}A_{t})x_{t}$
@@
Sampling and Generation.
During sampling (Algorithm 2), we integrate over the learned vector field to produce a new sample $x_{1}^{\prime}$. The learned vector field predicts Lie algebra elements that are integrated using Euler’s method with the exponential map $x_{t+\Delta t}=x_{t}\exp(\Delta tA_{t})$, ensuring that the trajectory remains on the Lie group manifold throughout generation. We also output the composed transform $M$, where $x_{1}^{\prime}=Mx_{0}$ as it is required for generating new group elements.
@@
Algorithm 3 generates group elements $h$ consistent with the target group $H$. Given a data sample $x_{1}$, we transform it by $g\sim p(G)$ to obtain $x_{0}=gx_{1}$. We then run Alg. 2 partially (lines 3-11) on $x_{0}$, obtaining outputs $x_{1}^{\prime}$ and $M$. The composed transform $Mg$ forms a group element $h\in H$ that maps $x_{1}$ to $\hat{x}_{1}$. Note that the new sample $x_{1}^{\prime}$ does not always match $x_{1}$ and in fact converges to the closest mode in the orbit (see Section 5.4 for more details), and thus this procedure does not always produce the identity element.
Comparison with Standard Flow Matching.
While standard flow matching operates in Euclidean space, our approach operates on Lie group manifolds, learning vector fields in the tangent space (Lie algebra) that generate curved trajectories respecting the group structure. The exponential and logarithm maps replace Euclidean interpolation with geodesic-like paths on the group manifold. More importantly, we learn a generative model of transformations rather than data points, yielding a lower-dimensional problem on the group manifold itself, not over the entire data space. We formulate symmetry discovery as flowing from a broad hypothesis group to the specific symmetries in the dataset, with challenges analyzed in Section 5.4.
@@ 5 Experiments
To evaluate LieFlow , we generate several datasets with known symmetries of simple 2D and 3D point clouds of canonical objects (Figure 9 in Appendix B). For the 2D datasets, we consider the target groups $H=C_{4},D_{4}$ and perform flow matching from $\mathrm{SO}(2)\rightarrow C_{4}$, $\mathrm{GL}(2,\mathbb{R})^{+}\rightarrow C_{4}$, and $\mathrm{GL}(2,\mathbb{C})\rightarrow D_{4}$. To learn $D_{4}$ symmetries, we perform flow matching over the complex domain, where the data samples are first converted into complex numbers and passed through the complex-valued network, which uses twice as many neurons. We find that $20$ steps for inference works well. For the 3D datasets, we use an object with no self-symmetries and consider the tetrahedral $\mathrm{Tet}$, octahedral $\mathrm{Oct}$ and rotation over $z$ axis $\mathrm{SO}(2)$ as target groups. We consider flow matching from $\mathrm{SO}(3)\rightarrow\mathrm{Tet}$, $\mathrm{SO}(3)\rightarrow\mathrm{Oct}$, $\mathrm{SO}(3)\rightarrow\mathrm{SO}(2)$ and use $100$ steps at inference. To quantify and comapre the quality of the learned symmetries, we compute the Wasserstein-1 distance between the generated group elements and the ground truth group elements. For all experiments, we use a simple 3-layer MLP with GELU (hendrycks2016gaussian) activations. See Appendix B for more details.
5.1 2D Datasets
(a) $\mathrm{SO}(2)$ to $C_{4}$
(b) $\mathrm{GL}(2,\mathbb{R})^{+}$ to $C_{4}$
(c) $\mathrm{GL}(2,\mathbb{C})$ to $D_{4}$
Figure 2: 2D Datasets: Generated data samples. The samples match closely to the original dataset symmetries of $C_{4}$ or $D_{4}$.
Figure 2 shows generated samples from random $x_{1}$ in the test set. We see that most samples correctly align with the target distributions $C_{4}$ or $D_{4}$. Surprisingly, we find that flow matching in the complex domain, from $GL(2,\mathbb{C})$ to $D_{4}$, correctly recovers all 8 group elements. Considering that we are flowing from continuous groups to discrete groups, our method works quite well, with little variance at the target modes.
Figure 3 visualizes the centroids of a $100$ generated samples over $t\in[0,1]$. The $x$ coordinate of the centroids are offset as $t$ increases. As expected, we can see the centroids converge to $4$ modes for the $C_{4}$ target groups and $8$ for the $D_{4}$ group as $t\to 1$. These figures show that the learned distribution over the prior group becomes quite peaky, suggesting that it does indeed learn a discrete group quite well and learns the correct modes. Specifically for $\mathrm{SO}(2)$, we can decompose the learned transformation matrices (which should be very close to elements of $C_{4}$) and plot a histogram of the angles. As we have the ground truth transforms of the dataset, we first canonicalize $M$ in Alg. 2 and perform polar decomposition over it to obtain the rotation matrix and finally convert them to angles. Figure 3(d) shows that the learned distribution gives the correct peaks at the $C_{4}$ elements and is close to the correct distribution (mixture of 4 delta distributions).
(a) $\mathrm{SO}(2)$ to $C_{4}$
(b) $\mathrm{GL}(2,\mathbb{R})^{+}$ to $C_{4}$
(c) $\mathrm{GL}(2,\mathbb{C})$ to $D_{4}$
(d) Angles
Figure 3: 2D Datasets: Visualization of trajectories from $t=0$ (left) to $t=1$ (right) of the centroids of the transformed objects over $100$ samples, and a histogram of learned angles for $\mathrm{SO}(2)$ to $C_{4}$.
(a) $\mathrm{SO}(2)$ to $C_{4}$
(b) $\mathrm{GL}(2,\mathbb{C})$ to $D_{4}$
Figure 4: Time progression of $x_{t}$ when generating $5$ samples over 2020 steps. The gray arrow shows the original $x_{1}$ and the color represents the determinant of the generated transformation matrix.
Figures 4(a), 4(b) show the intermediate timesteps during sample generation for $\mathrm{SO}(2)\to C_{4}$ and $\mathrm{GL}(2,\mathbb{C})\to D_{4}$, respectively ($\mathrm{GL}(2,\mathbb{R})^{+}\to C_{4}$ is given in Appendix C.1). Interestingly, we see that depending on the random transformation to produce $x_{0}$ (line 3 in Alg. 2), the final generated sample does not match the original $x_{1}$ and seems to flow to the closest group element in the orbit of $x_{1}$ of the target group, i.e., $\operatorname*{arg\,min}_{gx_{1}}\{\|gx_{1}-x_{0}\|:g\in G\}$. As the dataset contains all $C_{4}$ or $D_{4}$-transformed samples of the object, the flow still produces samples that are close to the data distribution. Furthermore, the progression shows that the vector field $v_{t}$ is close to 0 for the earlier timesteps (the determinants are close to $1$) and only induces a flow towards the closest $gx_{1},g\in T$ when $t$ is close to $1$. We analyze this “last-minute mode convergence” phenomenon in Section 5.4.
5.2 3D Datasets
@@
(c) $\mathrm{SO}(3)$ to $\mathrm{Oct}$ (power schedule)
(d) $\mathrm{SO}(3)$ to $\mathrm{SO}(2)$
@@
Figure 5: 3D Datasets: Visualization of 5,0005{,}000 generated elements of $\mathrm{SO}(3)$ by converting them to Euler angles. The first two angles are represented spatially on the sphere using Mollweide projection and the color represents the third angle. The elements are canonicalized by the original random transformation and the ground truth elements of the target group are shown in circles with black borders. The points are jittered with uniformly random noise to prevent overlapping.
Figure 5 visualizes generated samples after training as points on $\mathrm{SO}(3)$ and the ground truth symmetries are shown with transparent circles with black borders. For the tetrahedral group (Figure 5(a)), our model learns to correctly output transformations close to $\mathrm{Tet}$ elements, with some errors. For the octahedral group (Figure 5(b)), our flow model is unable to extract the correct symmetries and learns virtually a uniformly random distribution over the prior $\mathrm{SO}(3)$. For $\mathrm{SO}(2)$ (Figure 5(d)), our model correctly learns to output rotations around the $z$-axis, with some small errors. Figure 12 (Appendix C.2) shows visualizations analogous to Figure 3, where our model seems to learn $\mathrm{Tet}$ and $\mathrm{SO}(2)$ elements, creating some visible clusters, but fails to discover $\mathrm{Oct}$ symmetries.
We also generate transformed point clouds given by our FM model for $\mathrm{Tet}$ and $\mathrm{SO}(2)$, as shown in Figure 13 (Appendix C.2). We can see several repeated point clouds, and see that some samples are rotations or reflections of other samples. We also show the intermediate $x_{t}$ during generation in Figure 16 and Figure 17 (Appendix C.2). For $\mathrm{Tet}$ group, as in the 2D case, we see that the transformed point clouds converge to the closest group element in the orbit. For the $\mathrm{SO}(2)$ group, we can see that the model produces rotations to turn one node of the tetrahedron to be aligned with the $z$-axis and one triangle face to be in the $xy$-plane, which corresponding to $\mathrm{SO}(2)$ symmetries around the $z$-axis. For the $\mathrm{Oct}$ group, we show its generated point clouds and the intermediate $x_{t}$ in Appendix C.2, where we can see that the FM model produces outputs close to the identity at most steps and fails to properly discover $\mathrm{Oct}$.
Motivated by our analysis on “last-minute mode convergence” in the next subsection, where we observe that the entropy of the posterior remains stationary until large times, we introduce a novel time scheduling scheme to sample more timesteps near $t=1$. Instead of sampling time from a uniform distribution $\mathcal{U}(0,1)$, we propose using a power distribution where the density is computed as $p(x)=nx^{n-1},x\sim\mathcal{U}(0,1)$ and $n$ is the skewness parameter. The effect of different skewness values on the density is shown in Figure 19 (Appendix C.3). In our experiments, we chose $n=5$.
In Figure 5(c), our modified time schedule now enables our model, whose network is unchanged from before, to correctly learn the octahedral group. Figure 21(a) (Appendix C.3) shows that our model correctly distributes probability mass towards the $24$ $\mathrm{Oct}$ group elements. To test the limits of our time schedule, we also consider the icosahedral group, with $60$ elements. However, our method fails in this case, where the model seems to output the identity for all inputs (see Figure 21(b) in Appendix C.3). It is clear that while our time schedule can alleviate the last-minute mode convergence issue and allow our method to discover more complex groups, it still cannot handle more “symmetric groups”, i.e., higher-order subgroups.
5.3 Quantifying discovered symmetries
To quantify the quality of the discovered symmetries, we compute the Wasserstein-1 distance between the empirical distributions of the generated group elements and the ground truth group elements in the test set consist of 5,0005{,}000 samples. We choose LieGAN (yang2023generative) as our baseline since it is a recent generative approach that explicitly learns Lie algebra parameters for group-like transformations. For 2D datasets, we set the number of Lie generators to be 1, 2, and 4. For 3D datasets, we set the number of Lie generators to be 1, 3, and 6. The results are shown in Table 1 and Table 2. We can see that our method outperforms LieGAN in all experiments, demonstrating the effectiveness of our flow matching approach for symmetry discovery.

### v2

@@ 4.2 Training and Generation with LieFlow
8:  Take gradient descent step on $\nabla_{\theta}\left\|v_{\theta}(x_{t},t)-A\right\|^{2}$
9:until converged
Training.
Using the interpolation paths and the Lie group flow matching objective, we can now derive the training and sampling algorithms for our method. The training algorithm (Algorithm 1) directly implements the objective by sampling $x_{1}$ from the data distribution, sampling a transformation $g\sim p(G)$, and constructing $x_{0}=gx_{1}$. The target Lie algebra group element $A$ can be computed analytically using the sampled group element $g$. We sample a timestep $t\in[0,1)$ and construct a point $x_{t}=\exp(tA)x_{0}$ on the curve. Given $x_{t}$ and $t$, the model predicts $A$. We later show a modified time sampling procedure that works well for discovering discrete subgroups in Section 5.4.
Sampling and Generation.
During sampling (Algorithm 2), we integrate over the learned vector field to produce a new sample $x_{1}^{\prime}$. Starting from $x_{0}=gx_{1}$ for a random $g\sim p(G)$, we iteratively apply the predicted Lie algebra elements using Euler integration with the exponential map $x_{t+\Delta t}^{\prime}=\exp(\Delta_{t}A_{t})x_{t}^{\prime}$. This ensures that the trajectory remains on the group orbit.
@@
4:$M=I$
[…]
6:for $t=\{0,\Delta_{t},2\Delta_{t},\dots,(T-1)\Delta_{t}\}$ do
7:  $A_{t}=v^{\theta}(x_{t}^{\prime},t)$
8:  $x_{t+\Delta_{t}}^{\prime}=\exp(\Delta_{t}A_{t})x_{t}^{\prime}$
@@
We also accumulate the product of all applied transformations into $M$ such that at the end of sampling, $x_{1}^{\prime}=Mx_{0}^{\prime}$. The transformation $M$ is used for generating new group elements (Algorithm 3).
Algorithm 3 shows how to generate group elements $h$ consistent with the target group $H$. Given a data sample $x_{1}$, we transform it by $g\sim p(G)$ to obtain $x_{0}=gx_{1}$. We then run Alg. 2 partially (lines 3-10) on $x_{0}$, obtaining outputs $x_{1}^{\prime}$ and $M$. The key observation is that the composed transform MgMg forms a group element $h\in H$ that maps $x_{1}$ to $x_{1}^{\prime}$.
4.3 Handling Partially Observed Symmetries
The main assumption of this paper is that there exists a group $H$ to which the data distribution is invariant. This requires that the data be spread uniformly along the orbits of an entire group. In practice, however, data distributions can exhibit partial symmetries, i.e. $q(hx)=q(x)$ holds for only some $x\in\mathcal{X}$ and $h\in H$. We discuss how our method can be generalized to handle this case.
Formally, we consider the scenario where there exists a maximal subset $H_{x}\subseteq H$ for each $x\in\mathcal{X}$ such that for all $h\in H_{x}$, $q(hx)=q(x)$. For example, consider a data distribution of three arrow shapes: $q(\leftarrow)=q(\uparrow)=q(\rightarrow)=1/3$, and the group $H=\langle h\rangle$, where $h$ acts on data by clockwise $\pi/2$ rotation. Then we have $H_{\leftarrow}=\{e,h,h^{2}\}$, $H_{\uparrow}=\{e,h,h^{3}\}$, $H_{\rightarrow}=\{e,h^{2},h^{3}\}$. The goal of symmetry discovery then becomes, in addition to finding a subgroup $H\subset G$, identifying the input-dependent “symmetry subset” $H_{x}\subseteq H$ for any $x$.
Note that our method does not learn the explicit algebraic structure of $H$, but instead learns a distribution over the hypothesis group $G$ and identifies $H$ by (the closure of) the support of the learned distribution density. Analogously, to discover the partial symmetries stated above, it suffices to learn a conditional distribution over $G$, $p_{\theta}(\cdot|x)$, whose support at each $x$ reveals the symmetry subset $H_{x}$.
In Algorithm 3, we first sample $x_{1}$ from the data distribution $q$ and then sample $h=Mg$ that maps $x_{1}$ to another point $x_{1}^{\prime}$ with an equal density. In other words, we can view $h$ as a sample from a conditional distribution $p_{\theta}(\cdot|x_{1})$. In our default setting of exact global symmetry, we do not need to specify the condition $x_{1}$ because the same symmetry group applies at every point. Algorithm 3 then simple performs ancestral sampling from the marginal $p_{\theta}(\cdot)=\int_{\mathcal{X}}p_{\theta}(\cdot|x_{1})q(x_{1})dx_{1}$. In the setting with partial symmetries, the same procedure instead queries an input-dependent symmetry subset $H_{x_{1}}\subset H$, revealed by the support of $p_{\theta}(\cdot|x_{1})$ at a given $x_{1}$. In Section 5.3, we show that our method can still discover the support of the underlying symmetry group even when the data only demonstrates partial symmetries.
(a) $\mathrm{SO}(2)$ to $C_{4}$
(b) $\mathrm{GL}(2,\mathbb{R})^{+}$ to $C_{4}$ (Lie algebra coefficients based sampling)
(c) $\mathrm{GL}(2,\mathbb{R})^{+}$ to $C_{4}$ (matrix composition based sampling)
(d) $\mathrm{SO}(2)$ to partial $C_{4}$
Figure 2: Histograms of $5,000$ learned angles for 2D datasets, where the peaks correspond to the elements of the target symmetry group $C_{4}$. LieFlow successfully recovers $C_{4}$ symmetry group from (a) a compact hypothesis group $\mathrm{SO}(2)$ (b) a non-compact hypothesis group, (c) different sampling strategies, and (d) even when the full symmetry group is partially observed. The peaks correspond to the elements of the target symmetry group $C_{4}$.
@@ 5 Experiments
To evaluate LieFlow, we consider a variety of datasets consisting of 2D and 3D point clouds with different underlying ground truth symmetry groups. We also show our method can succeed with different hypothesis groups, in settings with partially observed symmetries, where only a subset of group transformations are observed in data, and in settings where observations are noisy.
5.1 Datasets
We evaluate LieFlow on both synthetic 2D and 3D point clouds and on ModelNet10 (Wu et al., 2015), which contains 3D models of real-world objects. For both the synthetic and real-world objects, we generate several datasets by transforming the canonical objects with group elements sampled from a target ground truth symmetry group $H$, see Figures 56 in Appendix B.
Synthetic datasets
We consider two datasets, each generated from a single canonical object: a 2D arrow and a 3D irregular tetrahedron with no self-symmetries. For the 2D arrow, we add data points given by $H=C_{4}$ transformations and perform flow matching with hypothesis groups $\mathrm{SO}(2)$ and $\mathrm{GL}(2,\mathbb{R})^{+}$. For $\mathrm{SO}(2)$, we uniformly sample angles in $0,2\pi)$ to generate group elements. For $\mathrm{GL}(2,\mathbb{R})^{+}$, which is non-compact, we consider two different prior distributions to sample group elements $g$: (1) Lie algebra coefficients based sampling (L): uniformly sampling the coefficients of the four orthogonal basis vectors and applying the exponential map; (2) Matrix composition based sampling (M): independently sampling rotation, scaling, and shear matrices and multiplying them. See Appendix C for more details. For the 3D dataset, we attempt to discover four different ground truth symmetry groups: tetrahedral $\mathrm{Tet}$, octahedral $\mathrm{Oct}$, icosahedral $\mathrm{Ico}$ and $SO(2)$. We use $\mathrm{SO}(3)$ as a hypothesis group.
ModelNet10
To demonstrate that LieFlow can handle complex real-world objects of diverse shapes, we consider ModelNet10 dataset(Wu et al., 2015). We first randomly select $10$ objects from each of the ten categories of ModelNet10 as canonical shapes, and downsample the point clouds to $128$ points. See Figure 6 in Appendix B for example samples. We consider five different target symmetry groups $\mathrm{Tet}$, $\mathrm{Oct}$, $\mathrm{Ico}$, the $\mathrm{SO}(2)$ subgroup stabilizing the $z$ axis and $\mathrm{SO}(2)$ stabilizing $(0,\frac{1}{2},-\frac{\sqrt{3}}{2})$.
5.2 Evaluation
In order to properly measure the quality of discovered symmetries, we propose using the Wasserstein-1 distance between the generated group elements (sampled via Algorithm 3) and the ground truth elements. Given two distributions $P$ and $Q$ over a metric space $(\mathcal{M},d_{G})$, let $\Pi(P,Q)$ be the set of all couplings between $P$ and $Q$, the Wasserstein-1 distance is defined as:
$W_{1}(P,Q)=\inf_{\gamma\in\Pi(P,Q)}\mathbb{E}_{(x,y)\sim\gamma}[d_{G}(x,y)],$ (6)
and it can be computed using the POT library in Python
For all experiments, we use a simple five-layer MLP with RELU activations. As a baseline, we compare against LieGAN (Yang et al., 2023), a recent state-of-the-art method for learning Lie group symmetries using a GAN. We set the number of generators in LieGAN to be the same dimension as our hypothesis group, and keep the architecture of the discriminator the same as our predictive network for a fair comparison. See Appendix C for more details.
5.3 Results on Synthetic 2D Datasets
Figures 2(a), 2(b), and 2(c) show the histograms of learned angles for the 2D dataset with target group $C_{4}$. All histograms correctly display four peaks at the $C_{4}$ group elements, confirming that LieFlow discovers the underlying symmetries across different hypothesis groups and prior distributions.
Table 1: Wasserstein-1 distance between generated group elements and ground truth group elements in 2D experiments (lower is better). We use 11 and $4$ LieGAN generators, corresponding to the dimension of $\mathrm{SO}(2)$ and $\mathrm{GL}(2,\mathbb{R})^{+}$, respectively. (L) and (M) indicate Lie algebra coefficients based sampling and matrix composition based sampling, respectively.
Method $\mathrm{SO}(2)\to C_{4}$ $\mathrm{GL}(2,\mathbb{R})^{+}\to C_{4}$(L) $\mathrm{GL}(2,\mathbb{R})^{+}\to C_{4}$(M)
LieGAN (1.gen) $1.688$ $1.688$ $1.688$
LieGAN (4.gen) $1.672$ $1.672$ $1.672$
LieFlow(Ours) $\mathbf{0.072}$ $\mathbf{1.321}$ $\mathbf{0.428}$
Table 1 shows the Wasserstein-1 distance between the generated group elements and the ground truth group elements. LieFlow outperforms LieGAN in all settings. We can see that using a smaller hypothesis group $\mathrm{SO}(2)$ leads to much better performance, matching our intuition that a more informative prior is beneficial. Even with the larger $\mathrm{GL}(2)$ hypothesis group, both sampling settings outperform LieGAN, although the matrix composition prior (M) is superior.
@@ 5.3 Results on Synthetic 2D Datasets
(c) $\mathrm{SO}(3)$ to $\mathrm{SO}(2)$ over z axis
(d) $\mathrm{SO}(3)$ to $\mathrm{SO}(2)$ over $(0,\frac{1}{2},-\frac{\sqrt{3}}{2})$ axis
(e) $\mathrm{SO}(3)$ to $\mathrm{Ico}$, without power schedule
(f) $\mathrm{SO}(3)$ to $\mathrm{Ico}$, with power schedule
@@
Figure 3: 3D Datasets: Visualization of 5,0005{,}000 generated elements of $\mathrm{SO}(3)$ by Euler angles. The first two angles are represented spatially on the sphere using Mollweide projection and the color represents the third angle. The elements are canonicalized by the original random transformation and the ground truth elements of the target group are shown in circles with gray borders. The points are jittered with uniformly random noise to prevent overlapping. Our method recovers a variety of ground truth symmetry groups.
Partially Observed Symmetries. In this experiment, we consider the setting where only a subset of the transformations of the target group $C_{4}$ are present in the dataset, i.e. $C_{4}$ is only partially observed. Specifically, we construct the dataset by considering only $0,\frac{\pi}{2},\frac{3\pi}{2}$ degree rotations to the canonical arrow object with equal probability.
The results are shown in Figure 2(d). Even when only three transformations are observed in the dataset, LieFlow using Algorithm 3 can correctly recover all four elements of the group closure $C_{4}$.
5.4 Results on Synthetic 3D Datasets
Figures 3 visualize generated group elements for the different target symmetry groups on $\mathrm{SO}(3)$, where the points are generated from our flow model and the ground truth symmetries are shown as circles with gray borders. For the tetrahedral group (Figure 3(a)), octahedral group (Figure 3(b)), $\mathrm{SO}(2)$ group around the $z$-axis (Figure 3(c)), and $\mathrm{SO}(2)$ group around $(0,\frac{1}{2},-\frac{\sqrt{3}}{2})$ (Figure 3(d)), we can see that our model learns to correctly output transformations very close to the target group elements, similar to the 2D case. We find that training with a uniform time sampling procedure in Algorithm 1 fails to recover the icosahedral group and produces almost uniformly distributed elements over $\mathrm{SO}(3)$. We provide an improved sampling method in the next section.
(a) $\mathrm{SO}(3)$ to $\mathrm{Tet}$
(b) $\mathrm{SO}(3)$ to $\mathrm{SO}(2)$
Figure 4: Time progression of $x_{t}$ when generating $5$ samples over $10$ steps to illustrate the inference process. The object starts moving at later time steps in the discrete group case ((a)), whereas it gradually moves at each time step in the continuous group case ((b)). The gray object shows the original $x_{1}$.
Table 2(a) presents quantitative results, where LieFlow significantly outperforms LieGAN in all discrete group while achieving comparable performance on continuous groups.
Improved Time Sampling. To further analyze the failure of uniform time sampling, we first plot the time progression of transformed objects in Figure 4(a) and Figure 4(b). For discrete groups, the model only begins moving objects towards the group elements when $t$ is close to $1$, whereas for continuous groups, the transformation occurs gradually throughout the flow. We find that the entropy of the posterior remains stationary until $t$ approaches $1$ (see Appendix D for more details). This phenomenon of “last-minute convergence” makes learning discrete groups particularly challenging as the model needs to learn to produce a near-zero vector field for most time steps and only moves towards the correct mode at the end of inference.